# Ultrafast optical switching to a heterochiral charge-density wave state

### v1

@@ Discussion
The experimental evidence strongly suggests that the observed heterochiral CDW state is one of the manifestations of the 1T-TaS2 hidden state. This view is supported by the close correspondence of experimental conditions to prepare the heterochiral and hidden states, including the sample temperature, excitation parameters and fluence thresholds, electronic enhancement, surface quality, as well as the long metastable state lifetime and the ability to reverse the state generation.
Local switching of CDW chirality was previously reported for room-temperature 1T-TaS2 by an optical quench Zong2018 ; Vogelgesang2018thesis and low-temperature 1T-NbSe2 by voltage pulses Song2022 ; Liu2023 . In the room-temperature case, chirality switching is likely accomplished by passing through the intermediate, non-chiral incommensurate (IC) CDW phase Vogelgesang2018 ; Vogelgesang2018thesis ; Storeck2020 ; Storeck2021 . In our case of low-temperature 1T-TaS2, however, the fluences applied are far below those required for transiently reaching the IC phase, and a different, electronically driven mechanism must govern the switching, as further discussed below.
A qualitative description of the physical mechanism underlying the formation of the heterochiral CDW state must be able to explain the generation of a disordered state composed of translational and chiral domains through electronic excitation. The Kibble-Zurek mechanism generally predicts the formation of a high density of topological defects arising from a rapid quench through a second-order phase transition Deutschlander2015 ; Eggebrecht2017 . In our case, however, the system is not quenched from a high-symmetry to a broken-symmetry state, but rather from a symmetry-broken homogeneous phase into a texture of nearly degenerate domains with equally broken symmetries but different stackings or chiralities.
@@
The ultrashort electronic enhancement and vibrationally coherent character of the hidden-state preparation Maklar2023 ; Ravnik2018 provides a hint to the generation mechanism, which we discuss with the illustration in Fig. 3a. First, the electronic excitation by the laser pulse drastically deforms the free-energy potential and induces a ballistic structural motion towards an inverted CDW state. Such an inverted state is energetically highly unfavorable and was reported to exhibit large upward band shifts Zhang2022 ; Maklar2023 . Thus, we propose that the subsequent ballistic evolution of the system beyond the high-symmetry state ($\varphi=0$) and onto the high-energy barrier will cause a “forced symmetry breaking” (cf. Fig. 3a), as opposed to the spontaneous (Kibble-Zurek) symmetry breaking. As a result, the system locally branches off into many of the nearly degenerate configurations of the system. In this picture, the minority contribution of the $\beta$ chirality, as shown in our LEED images, can be attributed to its somewhat more unfavorable free energy and/or a narrower path across the transition state. The reversal to a monochiral state can also be understood on the basis of a shallow metastable energy minimum for a $\beta$ domain in an $\alpha$-domain environment.
Based on the present experimental evidence, we cannot assert with certainty the spatial distribution of generated $\beta$ domains. We carried out dynamical LEED simulations, which predict that only about 3.7 % of the total superstructure diffraction intensity stems from the subsurface layer (cf. Fig. S5c in the Supplementary Information). Therefore, in order to account for the observed 3-10 % of the total superstructure diffraction intensity contained in the $\beta$ peaks for different samples (Fig. 2b), a corresponding few-percentage minority of the surface layer (Fig. 3b, right) and/or a majority of the subsurface layer (Fig. 3b, left) will have to switch to the $\beta$ chirality. Recently, an STM study reported a state with mixed chiral domains after 1T-TaS2 crystals are exposed to a single laser pulse at a fluence of 3-5 mJ cm-${}^{\textrm{-}2}$ Ravnik2023 . This observation is in general agreement with the structural textures proposed here.
A full crystallographic reconstruction of the structure, as previously carried out for the homogeneous C-phase vonWitte2019 , is complicated by the intrinsic disorder of the hidden state. Specifically, we observe a reduced modulation of the energy-dependent LEED intensity (cf. Fig. S4b in the Supplementary Information) VanHove1986 , which likely follows from incoherent averaging over different stackings associated with translational nano-domains Gerasimenko2019 . Moreover, possible $\alpha$-$\beta$ cross peaks, which might be expected for a completely switched subsurface layer (Fig. 3b, left), are absent in our observations, which again suggests a surface layer of mixed $\alpha\beta$-domains.
@@
Irrespective of the specific details of the real-space texture, a general implication of our observations is the presence of extended interfaces of opposite CDW chirality at or near the surface, which causes the emergence of CDW moiré interfaces. In particular, the 27.8 rotational mismatch between CDWs in the $\alpha$ and $\beta$ layers gives rise to a commensurate CDW moiré superstructure (cf. Fig. S6b in the Supplementary Information) with an unusually large moiré length ($\lambda_{m}=43.68\,\text{\AA}$) compared to typical large-angle moiré materials Shallcross2010 ; Tilak2023 ; Carr2020 . Using density functional theory (DFT) simulations, we explore the electronic and structural order of the CDW moiré interface, considering a simplified atomic structure composed of two 1T-TaS2 trilayers of opposite CDW chirality (Fig. 4a, left). The unit cell of the superstructure contains 13 Ta hexagrams in each trilayer, in a fractal-type arrangement that forms a "super-hexagram" (see also Fig. S6b). The angle between $\alpha$ and $\beta$ super-hexagrams in adjacent layers creates a mixed stacking order Lee2019 (Fig. 4a, right). The periodically altering stacking is expected to lead to the collapse of interlayer dimerization Stahl2020 ; Ritschel2018 ; Lee2019 ; Petocchi2022 ; Nicholson2022 ; Salzmann2023 ; Lee2021 ; Butler2020 ; Fei2022 and locally varying conductivity, linking to the recent observation of the non-Drude-type conductivity of the hidden state Gao2022 .
@@
The main results of our DFT simulations are summarized in Fig. 4b-e. In the CDW moiré unit cell (cf. Fig. S6b), the relaxed atomic structure transforms from a ($\sqrt{13}\times\sqrt{13}$)R13.9 triangular superlattice to a $13\times 13$ kagome superlattice (Fig. 4b and Fig. S7), with the CDW chirality and Ta hexagrams of each layer remaining unaltered. The atomic displacements are of the order $0.04\,\text{\AA}$. They collectively induce an electronic reconstruction, with electron density redistributed into a novel double-ring charge texture (Fig. 4c, left) with a charge-depleted central hexagram. A prominent inner charge ring is composed of 6 hexagrams, while a less pronounced outer ring spans 12 hexagrams shared by adjacent unit cells. This texture may arise from the stacking-dependent electrostatic energy, as discussed in Sec. III of the Supplementary Information. Such an electron density distribution may correspond to the ring-like charge structures observed by STM in Refs. Ohta2021  and Ravnik2023 , hence favoring the $\alpha\beta$-domain texture (Fig. 3b, right). The regular $\beta$ domains observed in Ref. Ravnik2023 imply regions with a deeper heterochiral interface. We believe that future spectroscopic STM measurements will elucidate the connection between the observed ring-like charge pattern and the CDW moiré superstructure introduced here.
@@
Alongside the structural transformation and electronic reconstruction, the density-of-states calculation shows an emergent conducting peak at the Fermi energy ($E_{\textrm{F}}$), which strongly contrasts with the electronic gap of the C-phase (Fig. 4d and Fig. S9b). The band structure calculation further reveals multiple flat bands and Dirac cones emerging near $E_{\textrm{F}}$ (Fig. 4e). While the flat bands at $E_{\textrm{F}}$ may facilitate correlation effects Cao2018-1 ; Cao2018-2 , the lowest conduction bands (within 50 meV above $E_{\textrm{F}}$) feature characteristics of a kagome subsystem Yin2022 ; Dong2023 : a Dirac point at K, a Van Hove singularity at M and a flat band across $\Gamma$-K-M-$\Gamma$. The corresponding electron density distribution hosts a kagome network of conducting channels (Fig. 4c, right). It is worth noting that the nearly half-filled band at $E_{\textrm{F}}$ in Fig. 4d arises naturally from the heterochiral CDW lattice structure with the electronic reconstruction.
@@ Conclusions
Combining femtosecond laser excitation with high-coherence LEED, we observe a heterochiral CDW state at the surface of 1T-TaS2 crystals, featuring laser-controlled reversibility, ultrafast electronic enhancement, and a threshold behavior. Prepared under conditions corresponding to the well-known hidden state in this material, we suggest that coexistent CDW chiralities are a structural feature closely linked to this state. Furthermore, we propose that this new structural element arises from an electronically-induced forced symmetry breaking. The coexistent long-range order of opposite charge-density wave chiralities implies the presence of extended heterochiral interfaces, either between the first two surface trilayers or up to few layers below the surface. Using DFT simulations, we reveal the electronic reconstruction of the double-ring charge texture, identify an emergent kagome system, and obtain a nearly half-filled band at $E_{\textrm{F}}$. Future work may allow for large-area fabrication of CDW moiré interfaces. More broadly, our study demonstrates new possibilities for optical chiral engineering and light-induced moiré superstructures in quasi-two-dimensional materials.
Acknowledgements.

### v2

@@ Discussion
Figure 3: Proposed generation mechanism and possible structural texture for coexistent CDW 2D-chiralities. a, Suggested mechanism for the generation of a 2D-heterochiral hidden state with translational domains. The illustration depicts simplified free-energy surfaces as a function of the order parameter $\varphi$ of the homogeneous 2D-monochiral C phase, and along other coordinates of the high-dimensional configuration space (schematically indicated as multiple axes). The optically-induced displacive excitation of the CDW causes a rapid evolution of the system towards an inverted CDW state which is energetically highly unfavorable. Collision with this steep barrier leads to a forced symmetry breaking into a distribution of translational and 2D-chiral domains. Concurrent with electronic cooling and structural relaxation, this branching finally results in the formation of a metastable state with quenched disorder. b, From the relative diffraction intensities of $\alpha$ and $\beta$ peaks, the experiments imply a random distribution of both 2D-chiralities in the surface layer. Recent STM measurements corroborate this scenario [28]. Layers in both textures may be covered with translational domains, for which measured diffraction peak widths yield a minimum structural correlation length of about $\geq 20$ nm. This implies extended interfaces between CDWs of opposite 2D-chirality at or near the surface.
Figure 4: Emergent kagome system in the CDW moiré superstructure. a, Construction of the CDW moiré unit cell (see also depiction in Fig. S6b). (Left) The structure considered consists of two 1T-TaS2 trilayers of opposite CDW 2D-chirality, $\alpha$ (blue) and $\beta$ (red). (Right) The moiré structure is described by a super-hexagram unit cell containing 13 Ta hexagrams in each layer. Circles denoted by capital letters A to M indicate the center of the Ta hexagrams and the local stacking order of both layers [3], with darker filling color representing higher electrostatic energy. The inner and outer rings of the super-hexagram are marked by orange and purple, respectively. b, Kagome structural order emerging from the coupling of both layers. A comparison between atomic positions in the CDW moiré superstructure and the C-phase shows that displacements of the Ta hexagram centers (depicted by arrows) break the translational symmetry of Ta hexagram, resulting in a $13\times 13$ kagome superlattice. The concentric rings mark the inner and outer rings of the super-hexagrams. The Ta atoms at the center of the super-hexagrams experience no displacement. c, Energy- and band-resolved electron density distributions. The total electron density for bands from -$90$ meV to $60$ meV around $E_{\textrm{F}}$ develops a double-ring texture (left, red shading), while the kagome electronic subsystem (right, blue shading) features a conducting network (yellow arrows). Both charge orders follow the same kagome superlattice (black lines) as the atomic structure. Electron density of $\alpha$ and $\beta$ layers are shown in blue and red, respectively. Isosurface values are adjusted for optimal clarity of the charge pattern. d, Total density of states (DoS) and e, band structure of the 2D-heterochiral CDW state. Several flat bands and a kagome electronic subsystem (highlighted in blue) are visible near $E_{\textrm{F}}$.
.1 Nature and origin of the 2D-heterochiral CDW state
@@
The experimental evidence strongly suggests that the observed 2D-heterochiral CDW state is one of the manifestations of the 1T-TaS2 hidden state. This view is supported by the close correspondence of experimental conditions to prepare the 2D-heterochiral and hidden states, including the sample temperature, excitation parameters and fluence thresholds, electronic enhancement, surface quality, as well as the long metastable state lifetime and the ability to reverse the state generation.
Switching between different CDW orientations in 1T-TaS2 has first been observed in the NC-phase at room temperature with the switching involving a transient IC CDW state ($\varphi=0$°) [25, 27, 26], leading up to µm-sized domains . A transient IC state appears incompatible with the experimental fluences and equilibrium temperatures required for the creation of the hidden state. Therefore, the occurrence of both CDW 2D-chiralities in the hidden states requires a different mechanism, which will be discussed below. Moreover, Ravnik et al. [28] observed laser-induced mirrored domains across small areas via STM experiments, which exhibited a structural relaxation within minutes, and which were interpreted as a precursor to the formation of the H-state.
Pump-pump-probe experiments suggest that the hidden state creation depends on the hexagram amplitude mode, eventually involving a transient overshoot into an inverted hexagram state, i.e., the outer twelve Ta atoms are further away from the central Ta atom compared to the high-temperature $1\times 1$ phase [24]. THz spectroscopy and optical reflectivity measurements motivated a Landau-Ginzburg-Wilson theory to model the formation of the H-state after the optical pulse [9]. In this model, the H-state is described as consisting of domains of the 13 possible stackings (cf. Fig. 4a) with domain sizes around 10 nm [9]. Previous theoretical and experimental works have demonstrated that several of the 13 stackings lead to a metallic behavior [16, 30, 3, 35, 31, 32, 34, 33, 36]. A hidden state consisting of small domains appears consistent with a Drude-Smith-type conductivity of confined charge carriers as suggested by THz spectroscopy [9, 37]. .
@@
Based on the above discussion and results, the formation of the hidden state is qualitatively sketched in Fig. 3a. First, the electronic excitation by the laser pulse drastically deforms the free-energy potential by strongly populating the amplitude mode of the CDW, initiating an expansion to an inverted CDW state. The expansion from a CDW to an inverted CDW state requires passing close to the high-symmetry state ($\varphi=0$°). However, the inverted CDW state is energetically highly unfavorable, and from the high-symmetry state, the system can locally branch off into many of the nearly degenerate configurations, including different stackings and CDW 2D-chiralities (cf. Fig. 3b). In this picture, the minority contribution of the $\beta$ CDW orientation, as shown in our LEED images, can be attributed to its somewhat more unfavorable free energy and/or a narrower path across the transition state. The reversal to a 2D-monochiral state can also be understood on the basis of a shallow metastable energy minimum for a $\beta$ domain in an $\alpha$-domain environment.
This transition bears some resemblance to the Kibble-Zurek mechanism, which describes a quench through a second-order phase transition from a highly symmetric to a spontaneously broken symmetry state [38, 39]. In contrast, however, the H-state formation evolves from a symmetry-broken homogeneous phase after excitation into many symmetry-broken local stacking and mirror domains. This likely involves swiftly traversing the high-symmetry state, which motivates denoting the mechanism as a “forced symmetry breaking” compared to the spontaneous symmetry-breaking described by Kibble and Zurek.
Dynamical LEED simulations of 1T-TaS2 [1, 40] suggest that around 3.7% of the total signal arises from the subsurface S-Ta-S (tri-)layer (cf. Fig. S5, Supplementary Information). The signal of the $\beta$ peaks exceeds 5% for several samples, and reaches ten percent for two of the samples (cf. Fig. 2b). Therefore, it appears most likely that the $\beta$ domains are randomly distributed in both the surface and subsurface layers.
The rather large variation between the samples remains unclear, especially the two samples not showing signs of a plateau. Assuming that between 5 and 10% of a given layer is converted to a $\beta$ orientation and the domains being typically larger than 23 nm, up to 10-20% of a pair of layers is composed of an $\alpha$-$\beta$ interface. If the metallic conductivity of the hidden state arises from the local conductivities of the different domains, as discussed above based on Ref. [9], the $\alpha$-$\beta$ interface might affect the overall hidden state conductivity. Thus, in the following, we discuss the properties of the $\alpha$-$\beta$ interface based on density functional theory (DFT) simulations.
Properties of $\alpha$-$\beta$ interfaces
@@
The 27.8 rotational mismatch between CDWs between the $\alpha$ and $\beta$ layers gives rise to a commensurate CDW moiré superstructure (cf. Fig. S6b, Supplementary Information) with an unusually large moiré length ($\lambda_{m}=43.68\,\text{\AA }$) compared to typical large-angle moiré materials [41, 42, 43]. Using DFT simulations, we explore the electronic and structural order of the CDW moiré interface, considering a simplified atomic structure composed of two 1T-TaS2 trilayers of opposite CDW 2D-chirality (Fig. 4a, left). The unit cell of the superstructure contains 13 Ta hexagrams in each trilayer, in a fractal-type arrangement that forms a “super-hexagram” (see also Fig. S6b, Supplementary Information). The angle between $\alpha$ and $\beta$ super-hexagrams in adjacent layers creates a mixed stacking order [3] (Fig. 4a, right).
@@
The main results of our DFT simulations are summarized in Fig. 4b-e. In the CDW moiré unit cell (cf. Fig. S6b, Supplementary Information), the relaxed atomic structure transforms from a ($\sqrt{13}\times\sqrt{13}$)R13.9 triangular superlattice to a $13\times 13$ kagome superlattice (Fig. 4b and Fig. S7, Supplementary Information), with the CDW 2D-chirality and Ta hexagrams of each layer remaining unaltered. The atomic displacements are of the order $0.04\,\text{\AA }$. They collectively induce an electronic reconstruction, with electron density redistributed into a novel double-ring charge texture (Fig. 4c, left) with a charge-depleted central hexagram. A prominent inner charge ring is composed of 6 hexagrams, while a less pronounced outer ring spans 12 hexagrams shared by adjacent unit cells. This texture may arise from the stacking-dependent electrostatic energy, as discussed in Sec. III of the Supplementary Information. Such an electron density distribution may correspond to the ring-like charge structures observed by STM in Refs. [21] and [28], hence favoring the $\alpha\beta$-domain texture (Fig. 3b, right). The regular $\beta$ domains observed in Ref. [28] imply regions with a deeper 2D-heterochiral interface. We believe that future spectroscopic STM measurements will elucidate the connection between the observed ring-like charge pattern and the CDW moiré superstructure introduced here.
@@
Alongside the structural transformation and electronic reconstruction, the density-of-states calculation shows an emergent conducting peak at the Fermi energy ($E_{\textrm{F}}$), which strongly contrasts with the electronic gap of the C-phase (Fig. 4d and Fig. S9b, Supplementary Information). The band structure calculation further reveals multiple flat bands and Dirac cones emerging near $E_{\textrm{F}}$ (Fig. 4e). While the flat bands at $E_{\textrm{F}}$ may facilitate correlation effects [44, 45], the lowest conduction bands (within 50 meV above $E_{\textrm{F}}$) feature characteristics of a kagome subsystem [46]: a Dirac point at K, a van Hove singularity at M and a flat band across $\Gamma$-K-M-$\Gamma$. The corresponding electron density distribution hosts a kagome network of conducting channels (Fig. 4c, right). It is worth noting that the nearly half-filled band at $E_{\textrm{F}}$ in Fig. 4d arises naturally from the 2D-heterochiral CDW lattice structure with the electronic reconstruction.
The DFT simulations indicate that a kagome system emerges from the twisting of the $\sqrt{13}\times\sqrt{13}$R13.9° CDW of 1T-TaS2 by 27.8° with the hexagrams as large “virtual” atoms. The $\sqrt{13}\times\sqrt{13}$ CDW occurs in several transition metal dichalcogenides (TMDCs) and it would be interesting to study if the proposed emergent kagome system is specific to 1T-TaS2 or rather connected to the basic geometry, i.e., occurring in several related compounds. Ohta et al. [21] studied an $\alpha$-$\beta$ structure of 1T-TaSe2 via STM and found a ring structure resembling Fig. 4c, with a prominent inner ring of six hexagrams and a depleted signal at the center and in the outer ring. These related observations suggest the design of structures connecting the features of moiré and kagome systems, as also discussed in Ref. [46].
@@ Conclusions
Combining femtosecond laser excitation with high-coherence LEED, we observe the coexistence of both CDW orientations at the surface of 1T-TaS2 crystals. Prepared under conditions corresponding to the well-known hidden state in this material, we suggest that the coexistence of both CDW orientations is a structural feature closely linked to the hidden state. Furthermore, we propose that this new structural element arises from a forced symmetry breaking leading to domains with different stackings and CDW orientations. The coexistent ferro-rotational domains $\geq 20$ nm in diameter implies the presence of extended $\alpha$-$\beta$ interfaces. Using DFT simulations, we reveal the electronic reconstruction of the double-ring charge texture, identify an emergent kagome system, and obtain a nearly half-filled band at $E_{\textrm{F}}$. At present, it is uncertain to what degree the emergent metallicity of $\alpha$-$\beta$ interfaces contribute to the macroscopic properties of the hidden state, as this likely depends on the actual fraction of interfaces formed. However, future work will allow for large-area fabrication of such CDW moiré interfaces, enhancing twist-angle electronic engineering by 2D-chirality.
Methods
LEED measurements
[…]
Electron-laser spatial overlap
[…]
Generation of single- and double-pulse
The optical quench of 1T-TaS2 samples is conducted using either single or double laser pulses. Single pulses are isolated with a Pockels cell inside a laser amplifier (center wavelength 1030 nm). Femtosecond excitation pulses of about 200 fs duration (center wavelength 800 nm, FWHM spectral width 7 nm) are produced by an optical parametric amplifier. Double pulses with a pulse separation $|\tau|\leq 2\,\textrm{ps}$ are generated by passing a single pulse through a Michelson interferometer. A half-inch polarizing beamsplitter cube is used to split the pulse, and a quarter-wave plate is used in each arm of the interferometer to optimize the total laser power output. The two output pulses are orthogonally polarized with a measured interferometric visibility of 3 %.
Reversal from the heterochiral to the monochiral state
The duration of the annealing laser pulse train is controlled by a mechanical shutter that produces a 1 s top-hat function with rising and falling edges spanning about 10 ms. Given sufficient fluence, the annealing laser pulse train can completely reverse the 2D-heterochiral state to a monochiral state (cf. Fig. 2c). This procedure is used for initializing each measurement. Incomplete reversal was observed if a sharper drop of the pulse train envelope was used.
Density functional theory simulations
Density functional theory (DFT) simulations are performed for the CDW moiré superstructure using projector augmented wave (PAW) [6] potentials as implemented in the Vienna Ab-initio Simulation Package (VASP) [51, 4]. A 350 eV kinetic energy cutoff in the plane-wave expansion is used. The PAW pseudopotentials correspond to the valence-electron configuration 5d36s2 for Ta and 3s23p4 for S. The exchange-correlation is treated within the generalized gradient approximation, as parameterized by Perdew, Burke, and Ernzerhof [8]. The Coulomb correlations within the 5d shells of Ta are described using the spherically averaged DFT+$U$ method [54], where the Hamiltonian only depends on the difference between on-site interaction $U$ and Hund’s exchange $J$, i.e., $U_{\textrm{eff}}=U-J$. We adopt the value of $U_{\textrm{eff}}=2.0$ eV for the Ta 5d orbitals [7]. Spin-orbit coupling is ignored in all simulations [30]. For structural relaxation, the lattice constants $a=43.68\,\text{\AA }$ (13 times of the $1\times 1$ primitive cell lattice constant $3.36\,\text{\AA }$) and $c=11.8\,\text{\AA }$ (twice the Ta-Ta interlayer distance $5.9\,\text{\AA }$) are fixed [1, 56, 57], while the positions of 1014 atoms are relaxed by a $\Gamma$-centered $1\times 1\times 2$ k-point mesh until the maximum force on each atom becomes less than 10 meV/Å. A $\Gamma$-centered $3\times 3\times 2$ k-point mesh is used for calculating the self-consistent ground state charge and electronic band structures. The total energies are converged to 10-$^{\textrm{-}4}$ eV. In a separate calculation, The 1T-TaS2 C-phase is simulated by a 156-atom AL-stacking unit cell using a $\Gamma$-centered $6\times 6\times 6$ k-point mesh [30, 3].